OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x – 2y = 8. (4 points) x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, –4) x intercept = (–8, 0); y intercept = (0, –4) x intercept = (–2, 0); y intercept = (0, 8)

OpenStudy (anonymous):

4. Jimmy's Chocolate Shop sells two different fudge bars. The dark chocolate fudge bars sell for $5 per pound and the milk chocolate fudge bars sell for $4 per pound. You purchased a combination of these bars for your family and friends and your bill came to $150. This scenario can be modeled by the equation 5d + 4m = 150. Which of the following is not a possible combination of dark chocolate fudge bars and milk chocolate fudge bars? (4 points) 6 dark bars and 30 milk bars 10 dark bars and 25 milk bars 12 dark bars and 23 milk bars 14 dark bars and 20 milk bars 5. Identify the correct slope and y intercept of the equation x – 4y = 4. (4 points) slope = 4; y intercept at (0, –1) slope = ; y intercept at (0, 1) slope = ; y intercept at (0, –1) slope = –4; y intercept at (0, 1)

OpenStudy (anonymous):

the answer to the first question is the second choice

OpenStudy (anonymous):

4. 5d + 4m = 150 process of elimination... a.) 5(6) + 4(30) = 150 30 + 120 = 150 not the answer b.) 5(10) + 4(25) = 150 50 + 100 = 150 not the answer c.) 5(12) + 4(23) = 150 60 + 92 = 152 ------> THE ANSWER d.) 5(14) + 4(20) = 150 70 + 80 = 150 not the answer

OpenStudy (anonymous):

To find x int- y=0 To find y int- x=0 x-2y=8 (0)-2y=8 y=-4 so y int is 0,-4 x-2y=8 x-2(0)=8 x=8, x int is 8,0

OpenStudy (anonymous):

@ThatHippieKid try posting one question per post